There is a row of cars starting a motor car race.If the 3rd car after the middle car is in the 10th position.How many cars are there at the starting?
1
Expert's answer
2019-08-12T09:02:38-0400

If there is the middle car we can find total number of cars as


"N=2k+1"

The position of the middle car is


"k+1"

The position of the 3rd car after the middle car is 


"k+1+3=10"

"k=10-4=6"

So,


"N=2(6)+1=13"